# Insights from leptohadronic modelling of the brightest blazar flare

### v1

@@ 4.5 Contribution of the 4FGL FSRQs to the IceCube neutrino flux at energies $E^{o}_{\nu}\geq 100$ TeV
The estimated average neutrino yield in IceCube from all 4FGL FSRQs is $\sim 6\times 10^{-2}$ muon neutrinos per year given the linear relation between the $\nu$ and $\gamma$-ray integral fluxes and Eq. (41) for linking neutrino SEDs and Fermi-LAT light curves from the LCR. Given $\approx 12$ “pure” muon neutrinos per year observed by IceCube, it implies the 4FGL FSRQs contribute at the level up to $\sim 0.5$ of the IceCat-1 yearly muon neutrino rate at energies $E^{o}_{\nu}\geq 100$ TeV.
@@ 4.5.1 The case of TXS 0506+056
Although TXS 0506+056 is not classified as an FSRQ in the 4FGL catalogue (Abdollahi et al., 2020), it still may belong to this class as pointed out by Padovani et al. (2019) showing that this source may host BLR and DT external photon fields and a model similar to the one used for 3C 454.3 may describe the geometry of this source. We apply the procedure described in Sects. 3.4 and 3.5 to TXS 0506+056. This results in $\sim 4\times 10^{-4}$ IceCube muon neutrinos per year on long-term average and in $\sim 10^{-3}$ muon neutrinos per year within ±$\pm 180$ days around the IceCube-170922A arrival time (at energies $E^{o}_{\nu}\geq 100$ TeV) which lies within the range of different model predictions of Keivani et al. (2018, table 9) who performed dedicated modelling for this source specifically with various parametrization of the external photon field.
@@ 5.1 Location of the $\gamma$-ray emitting region beyond $R_{\mathrm{BLR}}$
The location of the dissipation radius beyond the outer radius of the BLR but not very far from it is in line with recent studies (see, however, Appendix E): Zheng et al. (2017) within leptonic models for 36 FSRQs obtained fits indicating the location of the emitting zone beyond the outer radius of the BLR; Costamante et al. (2018), after analyzing around seven years of Fermi-LAT data, did not find features in the $\gamma$-ray spectrum indicating presence of strong $\gamma\gamma$ absorption concluding that “$\gamma$-rays seem either produced outside the BLR most of the time, or the BLR is $\sim 100$ times larger than given by reverberation mapping”; Meyer et al. (2019) in a similar Fermi-LAT data analysis of a longer observation period concluded that the emitting region should be located beyond $1.36R_{\mathrm{BLR}}$ although their BLR model is somewhat different from the one in our work. Wang et al. (2022) suggested that multi-wavelength flares (contemporaneous optical, $X-$, and $\gamma$-ray) should occur at intermediate distances from the SMBH as in the case for the flare of 3C 454.3 we considered.
@@ 5.3 Contribution of Fermi-LAT FSRQs to the neutrino flux
Protons can have an acceleration timescale different from the one of electrons. By means of a simple scan (see Appendix D), we obtained that for $\xi_{\mathrm{acc}}\sim 5\times 10^{4}$ neutrino production reaches maximum and is $\sim 3$ times more efficient than for the case of $\xi_{\mathrm{acc}}\sim 10^{7}$. Thus, for a model with a separate neutrino production zone inside the BLR and optimal acceleration timescale, one can increase the neutrino yield by a factor of $\sim 2\times 10^{2}$ raising the potential contribution of 4FGL FSRQs to the IceCat-1 neutrinos up to $100$%. This is just a rough estimate and does not take into account possible changes in other model parameters. The consideration of a model with several emission zones is beyond the scope of the current paper.
@@ 5.4 A possible delay between neutrino emission and electromagnetic flares in blazars
For neutrino SEDs peaking at a few hundred TeV – a few PeV, the energy-loss time scale of proton photopion production exceeds the timescale of a typical multi-wavelength flare by almost three orders of magnitude. Depending on the Lorentz factor, dissipation radius and parameters of the external photon field, if the conditions for $p\gamma$ interactions hold for sufficiently long, the time window within which a significant neutrino fluence is expected can be as wide as $\sim 10^{3}$ d implying the possibility of a substantial months-to-years lag between a potential neutrino detection and the corresponding multi-wavelength flare. This circumstance can be accounted for in future searches of correlations between high-energy neutrinos and their electromagnetic counterparts.
@@ 6 Conclusions
If protons are co-accelerated with electrons, the X-ray data constrain the proton energy density to be at maximum $\rho_{p/e}\approx 1.3\times 10^{2}$ times greater than the electron energy density. By applying the same approach to four other periods of activity of 3C 454.3 we determined that the integral neutrino and $\gamma$-ray fluxes are approximately linearly proportional. Applying the obtained $\nu-\gamma$ scaling to the whole period of Fermi-LAT observations using the light curve from the Fermi LCR we constructed a neutrino light curve of 3C 454.3. We estimate that on average 3C 454.3 produces $6\times 10^{-3}$ IceCube muon neutrinos per year at energies $E^{o}_{\nu}\geq 100$ TeV in the framework of the single-zone “standing-feature” leptohadronic model. The proton maximum energy is not directly constrained by the electromagnetic flaring SED of 3C 454.3. Under more optimistic assumptions about the maximum proton energy (i.e. larger values $E^{\prime\mathrm{max}}_{p}$, lower values $\xi_{\mathrm{acc}}$), the maximum neutrino yield is similar or a factor of a few lower than in our benchmark model but the peak neutrino energy is approximately 100 PeV, consistent with the energy of the recently discovered $>$100 PeV neutrino KM3-230213A (The KM3NeT Collaboration, 2025).
Extrapolation of the 3C 454.3 model to the whole list of Fermi-LAT 4FGL FSRQs shows that they can account for up to $\sim 0.5$% of the IceCat-1 neutrino alerts at energies $E^{o}_{\nu}\geq 100$ TeV implying that FSRQs during multi-wavelength flares are not the main source of $\geq 100$ TeV neutrinos observed by IceCube. The total expected yield from Fermi-LAT FSRQs in the planned TRIDENT and IceCube-Gen2 neutrino telescopes will be about $\dot{N}_{\mathrm{yr}}(E^{o}_{\nu}\geq 100\mathrm{\,TeV})\sim 1\times(\rho_{p/e}/10^{2})$ muon neutrinos per year. Next-generation neutrino telescopes are thus expected to detect approximately one multimessenger ($\gamma$ + $\nu_{\mu}$) flare per year from FSRQs. We have shown that the electromagnetic flare peak can precede the neutrino arrival by months to years and that the Lorentz factor of the jet plays a crucial role in determining this time delay.
@@ Acknowledgements
We thank Karri Koljonen for help with the ISIS software, D. Ehlert for sharing his functions for calculating photon fields from the AD, and M. Kachelrieß, D. Khangulyan, T. Dzhatdoev, M. Turchetta, and J. Simpson for helpful discussions.
This work made use of data supplied by the UK Swift Science Data Centre at the University of Leicester. Part of this work is based on archival data, software or online services provided by the Space Science Data Center – ASI. This research made use of the Fermi-LAT data provided by NASA Goddard Space Flight Center. This paper has made use of up-to-date SMARTS optical/near-infrared light curves that are available at www.astro.yale.edu/smarts/glast/home.php. Data from the Steward Observatory spectropolarimetric monitoring project were used. This research has made use of the NASA/IPAC Infrared Science Archive, which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. This work made use of the following software packages: python (Van Rossum & Drake, 2009), astropy (Astropy Collaboration et al., 2013, 2018, 2022), Jupyter (Perez & Granger, 2007; Kluyver et al., 2016), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), scipy (Virtanen et al., 2020, 2021), astroquery (Ginsburg et al., 2019, 2024), pandas (The pandas development team, 2020; Wes McKinney, 2010), fermitools 2.2.0, fermipy (v. 1.2.2) (Wood et al., 2017), Xspec (Arnaud, 1996), iMINUIT (Dembinski et al., 2021; James & Roos, 1975), snakemake (Mölder et al., 2021), pyLCR (Abdollahi et al., 2023) [^18] https://github.com/dankocevski/pyLCR, numba (Lam et al., 2015, 2022), and tqdm (da Costa-Luis et al., 2023). This research has made use of NASA’s Astrophysics Data System and Pathfinder (Iyer et al., 2024). Software citation information aggregated using The Software Citation Station (Wagg & Broekgaarden, 2024a, b).
@@
All data used for obtaining the observed SEDs are publicly available at corresponding web resources of the instruments cited in Sect. 2. The AM3 program for modelling the observable SEDs is publicly-available (Klinger et al., 2023) [^19] https://gitlab.desy.de/am3/am3. Our codes for producing the figures will be published upon acceptance of this manuscript.
@@ Appendix A On the systematic uncertainty of the UV, optical, and IR spectral-energy distributions
In this appendix, we compare Steward and SMARTS apparent magnitudes of 3C 454.3 measured within the same nights over the whole history of their joint observations in the R and V bands. The corresponding scatter plots are shown in Figs. 13 and 14 respectively. The root-mean-square difference (i.e. the square root of the mean squared difference of the magnitude) and the mean difference between Steward and SMARTS magnitudes are displayed in the titles of Figs. 13 and 14. There is a systematic shift in the values of magnitudes between Steward and SMARTS ($0.02-0.04$) which is subdominant to the root-mean-square scatter ($0.08-0.09$). An investigation of the reasons behind these discrepancies is beyond the scope of the paper. We take into account the observed discrepancy as an additional $8\%$ relative systematic uncertainty added quadratically to all IR, UV, and optical SEDs as stated in Sect. 2.5.
@@
Appendix B Corner plot for the MCMC fit to the seven-day averaged SED (MJD 55517–55524)
The corner plot is shown in Fig. 15. The values of the parameters and their mutual correlations are in line with the qualitative description of the results presented in Sect. 4.1.
Figure 15: The corner plot (Foreman-Mackey, 2016) obtained from the MCMC fitting (Foreman-Mackey et al., 2013) of the seven-day-averaged SED (MJD 55517–55524) with a “standing-feature” leptonic model. $t^{\prime}_{\mathrm{sim}}$ here is the same as for the one-day-averaged SEDs with $t^{o}_{\mathrm{av}}=1$ d (see Eq. 4). The corner plot is obtained from an MCMC containing $6,000$ steps per walker with $3\times 8=24$ walkers after combining chains from all the walkers (discarding $1,000$ initial steps of each walker as a burn-in) and applying thinning of $10$. Dashed vertical lines in the one-dimensional marginal histograms represent $16\%$, $50\%$, and $84\%$ percentiles of the posterior marginal distributions. Contour lines in the 2D marginal histograms represent $11.8\%$, $39.3\%$, $67.5\%$, and $86.4\%$ 2D containment regions.
@@ Appendix C Spectral-energy distributions for four other periods
$L^{\prime}_{e}$ [erg/s] $6.5\times 10^{43}$ $2.3\times 10^{43}$ $1.5\times 10^{43}$ $8.2\times 10^{42}$ $2.2\times 10^{43}$
@@
$R^{\prime}_{b}$ $2.9\times 10^{16}$ $1.8\times 10^{16}$ $1.5\times 10^{16}$ $5.8\times 10^{15}$ $1.2\times 10^{16}$
$B^{\prime}$ [G] $1.0$ $9.6\times 10^{-1}$ $1.7$ $2.2$ $1.0$
$E^{\prime\mathrm{min}}_{e}$ [eV] $2.2\times 10^{8}$ $1.7\times 10^{8}$ $2.1\times 10^{8}$ $1.1\times 10^{8}$ $8.3\times 10^{7}$
$E^{\prime\mathrm{max}}_{e}$ [eV] $3.6\times 10^{9}$ $6.9\times 10^{9}$ $2.8\times 10^{9}$ $4.2\times 10^{9}$ $1.8\times 10^{9}$
$\Gamma$ $13$ $11$ $22$ $19$ $32$
$\chi^{2}_{\mathrm{red}}$ $1.85$ $2.13$ $1.38$ $0.79$ $0.97$
@@
Figure 16: SEDs averaged over MJD 54700–54701 (a), MJD 54827–54833 (b), MJD 55180–55181 (c), MJD 55471–55472 (d) fitted with the “standing-feature” leptonic model in the same manner as SEDs in Fig 8. The best-fit values of the model parameters are shown in Table 4.

### v2

@@ 6.2 Contribution of the 4FGL FSRQs to the IceCube neutrino flux at energies $E^{o}_{\nu}\geq 100$ TeV
The estimated average neutrino yield in IceCube from all 4FGL FSRQs is $\sim 6\times 10^{-2}$ muon neutrinos per year given the linear relation between the $\nu$ and $\gamma$-ray integral fluxes and Eq. (40) for linking neutrino SEDs and Fermi-LAT light curves from the LCR. Given $\approx 12$ “pure” muon neutrinos per year observed by IceCube, it implies the 4FGL FSRQs contribute at the level up to $\sim 0.5$ % of the IceCat-1 yearly muon neutrino rate at energies $E^{o}_{\nu}\geq 100$ TeV. Our model-dependent upper limit is in agreement with (i) the work by the IceCube Collaboration (Aartsen et al., 2017), where it was shown that $27\%$ ($50\%$) of the diffuse neutrino flux in the energy range between $\sim 10$ TeV and $\sim 2$ PeV assuming its power-law spectral index of 2.5-2.5 (2.2-2.2) can be produced by the Fermi-LAT blazars from the second AGN catalogue; (ii) the work by Murase et al. (2018), which predicts that the blazar flares contribute up to $1\%-10\%$ of the sub-PeV neutrino diffuse flux; (iii) the paper by Hooper et al. (2019), which constrained the Fermi-LAT FSRQs’ contribution to the neutrino flux to be at the level of $\lesssim 10\%$; (iv) the study by Yoshida et al. (2023, see their fig. 12), where, for the linear relation between the $\gamma$-ray and $\nu$ fluxes, the Fermi-LAT 4LAC FSRQs were found to contribute up to $\lesssim 100\%$ of the IceCube diffuse high-energy neutrino flux. Thus, our rough model-driven estimate of the maximum contribution of Fermi-LAT FSRQs to the IceCube neutrino flux is lower by one to two orders of magnitude than the upper limits obtained in the references above and agrees with the conclusion by Abbasi et al. (2023b) that a small fraction, $1\%$, of bright AGNs emits neutrinos that pass the IceCube alert criteria.
@@ 6.2.1 The case of TXS 0506+056
Although TXS 0506+056 is not classified as an FSRQ in the 4FGL catalogue (Abdollahi et al., 2020), it still may belong to this class as pointed out by Padovani et al. (2019) showing that this source may host BLR and DT external photon fields and a model similar to the one used for 3C 454.3 may describe the geometry of this source. We apply the procedure described in Sects. 5.1 and 5.2 to TXS 0506+056. This results in $\sim 4\times 10^{-4}$ IceCube muon neutrinos per year on long-term average and in $\sim 10^{-3}$ muon neutrinos per year within $\pm 180$ days around the IceCube-170922A arrival time (at energies $E^{o}_{\nu}\geq 100$ TeV) which lies within the range of different model predictions of Keivani et al. (2018, table 9) who performed dedicated modelling for this source specifically with various parametrization of the external photon field.
@@ 7.1 Location of the $\gamma$-ray emitting region beyond $R_{\mathrm{BLR}}$
In contrast to the results of Bonnoli et al. (2011); Hu et al. (2015); Sahakyan (2021) who also modelled the SED of 3C 454.3 in its flaring states, we could not obtain a fit with the dissipation radius $x<R_{\mathrm{BLR}}$ lying within the BLR due to the $\gamma\gamma$ absorption as explained in Sect. 3.1.5. On the other hand, our conclusion $x>R_{\mathrm{BLR}}$ that the dissipation radius lies beyond the outer radius of the BLR but not very far from it is in agreement with recent studies by Zheng et al. (2017); Costamante et al. (2018); Meyer et al. (2019); Kundu et al. (2025). Wang et al. (2022) suggested that multi-wavelength flares (contemporaneous optical, $X$-, and $\gamma$-ray) should occur at intermediate distances from the SMBH as in the case for the flare of 3C 454.3 we considered, however this may be not the case for all blazars: e.g., FSRQ PKS 1424-418 shows evidence of $\gamma\gamma$ absorption inside the BLR during flares (Agarwal et al., 2024). Different conclusions about the location $x$ of the $\gamma$-ray emitting region may be explained by different assumptions about the AD luminosity (see Table 5) and the BLR size and/or geometry (Abolmasov & Poutanen, 2017; Naddaf et al., 2021; Naddaf & Czerny, 2024; Zhou et al., 2025; Hopkins, 2025). A systematic study of the influence of the uncertainties of the BLR and AD parameters on the estimation of $x$ is beyond the scope of the current work.
@@ 7.3 Contribution of Fermi-LAT FSRQs to the neutrino flux
Protons can have an acceleration timescale different from the one of electrons. By means of a simple scan (see Appendix E), we obtained that for $10^{4}\lesssim\xi_{\mathrm{acc}}\lesssim 10^{5}$ neutrino production reaches maximum and is $\sim 3$ times more efficient than for the case of $\xi_{\mathrm{acc}}\sim 10^{7}$. Thus, for a model with a separate neutrino production zone inside the BLR and optimal acceleration timescale, one can increase the neutrino yield by a factor of $\sim 2\times 10^{2}$, raising the potential contribution of 4FGL FSRQs to the IceCat-1 neutrinos up to 100100%. This is just a rough estimate and does not take into account possible changes in other model parameters. The consideration of a model with several emission zones is beyond the scope of the current paper.
@@ 7.4 A possible delay between neutrino emission and electromagnetic flares in blazars
For neutrino SEDs peaking at a few hundred TeV – a few PeV, the energy-loss time scale of proton photopion production exceeds the timescale of a typical multi-wavelength flare by almost three orders of magnitude. Depending on the Doppler factor, dissipation radius and parameters of the external photon field, if the conditions for $p\gamma$ interactions hold for sufficiently long, the time window within which a significant neutrino fluence is expected can be as wide as $\sim 10^{3}$ d implying the possibility of a substantial months-to-years lag between a potential neutrino detection and the corresponding multi-wavelength flare. This circumstance can be accounted for in future searches for correlations between high-energy neutrinos and their electromagnetic counterparts.
@@ 8 Conclusions
If protons are co-accelerated with electrons, the X-ray data constrain the proton energy density to be at maximum $\rho_{p/e}\approx 1.3\times 10^{2}$ times greater than the electron energy density. By applying the same approach to four other periods of activity of 3C 454.3, we determined that the integral neutrino and $\gamma$-ray fluxes are approximately linearly proportional. Applying the obtained $\nu-\gamma$ scaling to the whole period of Fermi-LAT observations using the light curve from the Fermi LCR we constructed a neutrino light curve of 3C 454.3. We estimate that on average 3C 454.3 produces $6\times 10^{-3}$ IceCube muon neutrinos per year at energies $E^{o}_{\nu}\geq 100$ TeV in the framework of the single-zone “standing-feature” leptohadronic model. The proton maximum energy is not directly constrained by the electromagnetic flaring SED of 3C 454.3. Under more optimistic assumptions about the maximum proton energy (i.e. larger values $E^{\prime\mathrm{max}}_{p}$, lower values $\xi_{\mathrm{acc}}$), the maximum neutrino yield is similar or a factor of a few different from that in our benchmark model but the peak neutrino energy can reach $\sim 100$ PeV, consistent with the energy of the recently discovered >>100 PeV neutrino KM3-230213A (KM3NeT Collaboration, 2025).
Extrapolation of the 3C 454.3 model to the whole list of Fermi-LAT 4FGL FSRQs shows that they can account for up to $\sim[0.5(\rho_{p/e}/130)]$% of the IceCat-1 neutrino alerts at energies $E^{o}_{\nu}\geq 100$ TeV implying that FSRQs during multi-wavelength flares are not the main source of $\geq 100$ TeV neutrinos observed by IceCube. The total expected yield from Fermi-LAT FSRQs in the planned TRIDENT and IceCube-Gen2 neutrino telescopes will be about $\dot{N}_{\mathrm{yr}}(E^{o}_{\nu}\geq 100\mathrm{\,TeV})\sim 1\times(\rho_{p/e}/10^{2})$ muon neutrinos per year. Next-generation neutrino telescopes are thus expected to detect approximately one multimessenger ($\gamma$ + $\nu_{\mu}$) flare per year from FSRQs. We have shown that the electromagnetic flare peak can precede the neutrino arrival by months to years and that the Doppler factor of the jet plays a crucial role in determining this time delay.
@@ Acknowledgements
We thank Karri Koljonen for help with the ISIS software, D. Ehlert for sharing his functions for calculating photon fields from the AD, and M. Petropoulou, X. Rodrigues, M. KachelrieSS, D. Khangulyan, T. Dzhatdoev, M. Turchetta, J. Simpson, and B. Sen for helpful discussions. We thank the anonymous Referee for their comprehensive review, which allowed us to improve the quality of the manuscript.
This work made use of data supplied by the UK Swift Science Data Centre at the University of Leicester. Part of this work is based on archival data, software or online services provided by the Space Science Data Center – ASI. This research made use of the Fermi-LAT data provided by NASA Goddard Space Flight Center. This paper has made use of up-to-date SMARTS optical/near-infrared light curves that are available at www.astro.yale.edu/smarts/glast/home.php. Data from the Steward Observatory spectropolarimetric monitoring project were used. This research has made use of the NASA/IPAC Infrared Science Archive, which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. This work made use of the following software packages: python (Van Rossum & Drake, 2009), astropy (Astropy Collaboration et al., 2013, 2018, 2022), Jupyter (Perez & Granger, 2007; Kluyver et al., 2016), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), scipy (Virtanen et al., 2020, 2021), astroquery (Ginsburg et al., 2019, 2024), pandas (pandas development team, 2020; Wes McKinney, 2010), fermitools 2.2.0, fermipy (v. 1.2.2) (Wood et al., 2017), Xspec (Arnaud, 1996), iMINUIT (Dembinski et al., 2021; James & Roos, 1975), snakemake (Mölder et al., 2021), pyLCR (Abdollahi et al., 2023) [^22] https://github.com/dankocevski/pyLCR, numba (Lam et al., 2015, 2022), and tqdm (da Costa-Luis et al., 2023). This research has made use of NASA’s Astrophysics Data System and Pathfinder (Iyer et al., 2024). Software citation information aggregated using The Software Citation Station (Wagg & Broekgaarden, 2024a, b).
@@
All data used for obtaining the observed SEDs are publicly available at the corresponding web resources of the instruments cited in Sect. 2. The AM3 program for modelling the observable SEDs is publicly-available (Klinger et al., 2023) [^23] https://gitlab.desy.de/am3/am3. The derived observational SED data points and our modelling codes are available at the Zenodo repository (Podlesnyi & Oikonomou, 2025).
@@ Appendix A On the systematic uncertainty of the UV, optical, and IR spectral-energy distributions
In this appendix, we compare Steward and SMARTS apparent magnitudes of 3C 454.3 measured within the same nights over the whole history of their joint observations in the R and V bands. The corresponding scatter plots are shown in Figs. 16 and 17 respectively. The root-mean-square difference (i.e. the square root of the mean squared difference of the magnitude) and the mean difference between Steward and SMARTS magnitudes are displayed in the titles of Figs. 16 and 17. There is a systematic shift in the values of magnitudes between Steward and SMARTS (0.020.040.02-0.04), which is subdominant to the root-mean-square scatter (0.080.090.08-0.09). An investigation of the reasons behind these discrepancies is beyond the scope of the paper. We take into account the observed discrepancy as an additional $8\%$ relative systematic uncertainty added quadratically to all IR, UV, and optical SEDs as stated in Sect. 2.5.
@@
Appendix B On the deabsorbed X-ray SED and its uncertainty
To account for the possible photoelectric absorption of X rays along the LoS, the solar abundance is adopted from Wilms et al. (2000) and the photoionization cross-section table $\sigma_{\mathrm{photoel}}(E_{i})$ is from Verner et al. (1996). The initial guess for the hydrogen column density $\kappa$ is set to $\kappa=6.81\times 10^{20}$ cm-2 according to the average value calculated with the use of the web tool [^24] https://heasarc.gsfc.nasa.gov/cgi-bin/Tools/w3nh/w3nh.pl? within the $0.1^{\circ}$-cone with the centre at the position of 3C 454.3 on H I maps obtained by HI4PI Collaboration et al. (2016), although the column density is left free along with the normalization $N_{X}$, break energy $E_{\mathrm{br}}$, and the two spectral indices ($\Gamma_{1}$ and $\Gamma_{2}$) of the broken-power-law X-ray spectrum. Using fit, plot_unfold, and write_plot ISIS functions, we extract both the observed $\mathrm{SED}_{o}(E_{i})$ and the deabsorbed $\mathrm{SED}_{d}(E_{i})$ from the best-fit model and take into account the uncertainty related to the deabsorption correction in the following manner:
$\sigma_{\mathrm{photoel}}(E_{i})=-\frac{\ln k_{a}(E_{i})}{\kappa},$ (44)
where $\varsigma_{d}(E_{i})$ is the final uncertainty of the deabsorbed SED after taking into consideration the uncertainty of the hydrogen column density $\varsigma_{\kappa}$, and $\varsigma_{o}(E_{i})$ is the statistical uncertainty of the observed $\mathrm{SED}_{o}(E_{i})$ in the $i$th energy bin $E_{i}$. It is the deabsorbed $\mathrm{SED}_{d}(E_{i})$ and its total uncertainty $\varsigma_{d}(E_{i})$ which is used further in modelling the multi-wavelength electromagnetic emission of 3C 454.3 $\varsigma_{\kappa}$ is estimated using conf_loop ISIS function calculating 68% confidence interval for each of the free parameters after having obtained the best-fit with the absorbed broken-power-law model. After obtaining $\kappa=7.7^{\,+1.8}_{-2.4}\times 10^{20}$ cm-2 for MJD 55517–55524, we perform the same fit as described in this paragraph for each of the seven one-day-averaged consecutive SEDs allowing $\kappa$ to vary only within the 68% confidence interval obtained above. This constraint is required to ensure the convergence of the fits for each of the individual days, since otherwise some fits fail to converge. The obtained best-fit values for each of the individual one-day-averaged SEDs as well as for the seven-day-averaged SED are shown in Table 2. The inferred values of $\kappa$ agree with the ones obtained by HI4PI Collaboration et al. (2016).
@@ Appendix C Spectral-energy distributions for four other periods
$L^{\prime}_{e}$ [erg/s] $6.5\times 10^{43}$ $2.3\times 10^{43}$ $1.5\times 10^{43}$ $8.2\times 10^{42}$ $1.7\times 10^{43}$
@@
$R^{\prime}_{b}$ [cm] $2.9\times 10^{16}$ $1.8\times 10^{16}$ $1.5\times 10^{16}$ $5.8\times 10^{15}$ $1.1\times 10^{16}$
$B^{\prime}$ [G] $1.0$ $9.6\times 10^{-1}$ $1.7$ $2.2$ $1.2$
$E^{\prime\mathrm{min}}_{e}$ [eV] $2.2\times 10^{8}$ $1.7\times 10^{8}$ $2.1\times 10^{8}$ $1.1\times 10^{8}$ $7.0\times 10^{7}$
$E^{\prime\mathrm{max}}_{e}$ [eV] $3.6\times 10^{9}$ $6.9\times 10^{9}$ $2.8\times 10^{9}$ $4.2\times 10^{9}$ $1.6\times 10^{9}$
$\Gamma$ $13$ $11$ $22$ 1919 $35$
$\chi^{2}_{\mathrm{red}}$ $1.73$ $1.95$ $1.38$ $0.70$ $0.76$
@@
Figure 18: SEDs averaged over MJD 54700–54701 (a), MJD 54827–54833 (b), MJD 55180–55181 (c), MJD 55471–55472 (d), fitted with the leptohadronic model (cf. Fig 13). The values of the model parameters are shown in Table 7. The ratio of proton and electron energy densities is fixed to $\rho_{p/e}=1.3\times 10^{2}$.
Appendix D Linear relation between integral $\nu$ and $\gamma$ fluxes
Figure 19: Illustration of the linear relation between the integral all-flavour $\nu$ flux of particles with energies $E^{o}_{\nu}\geq 100$ GeV and $\gamma$-ray flux of photons with energies $E^{o}_{\gamma}\geq 100$ MeV. The colours of data points correspond to the colours of the periods highlighted accordingly in Fig. 15. The error bars for neutrino fluxes represent the numerical uncertainty $\sim 40$% of normalisation of spectra of secondaries in hadronic processes (Cerruti et al., 2024)
As mentioned in Sect. 5.1, we obtained electromagnetic and neutrino SEDs of 3C 454.3 covering four other periods of the source’s activity. These periods are highlighted as vertical lines in Fig. 15, and their SEDs are shown in Fig. 18. In Fig. 19, we plot the integral particle fluxes of neutrinos and $\gamma$-rays for the five considered periods. The fluxes turn out to be linearly related. We explain this linearity as follows. As can be seen from Table 7, despite the change of the integral $\gamma$-ray photon flux of 3C 454.3 by two orders of magnitude, the electron injection luminosity changes by less than one order of magnitude, and most of the $\gamma$-ray flux change is driven by the change of the Lorentz factor $\Gamma$ due to the very high Compton dominance of 3C 454.3. We find that both $\gamma$ and $\nu$ fluxes are approximately proportional to $\delta^{5}$. At energies $E^{o}_{\nu}\gtrsim 100$ TeV, neutrinos begin to be directly produced on the BLR photon field, and Fermi-LAT $\gamma$ rays are produced via the IC scattering on the same photon field and the DT radiation boosted in the blob rest frame. At energies $E^{o}_{\nu}<100$ TeV, we find that the main target for the neutrino production is the $\sim$ MeV $\gamma$-rays produced via external Compton scattering on the DT photon field (with some contribution of the direct radiation from the AD, see Fig. 7). We confirm this by turning off the IC process in the simulation and observing no neutrinos with energies below $100$ TeV, while switching off the synchrotron radiation almost does not affect the neutrino production (and neutrinos with energies $E^{o}_{\nu}\gtrsim 100$ TeV, as expected, are not affected by these switches because they are produced on the external BLR photon field). The dependence of $\nu$ flux on the fifth power of the Doppler factor (having fixed other parameters) is expected (Dermer et al., 2012) and can potentially explain the recently observed evidence that radio-bright AGNs, spatially associated with high-energy neutrinos, have their estimated Doppler factors systematically larger than all radio-bright AGNs (Plavin et al., 2025).